OpenStudy (anonymous):

Change each improper fraction to a mix numeral Question: 27/5 Please explain it to me...SLOWY!! Thanks!

OpenStudy (saifoo.khan):

\[5\frac{2}{5}\]

OpenStudy (anonymous):

can u explain that to me?

OpenStudy (anonymous):

take alook

OpenStudy (anonymous):

OMG! tysm!!!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

i get it now ty!

OpenStudy (anonymous):

27/5= (25+2)/5 25/5+2/5= 5+2/5

OpenStudy (anonymous):

ok thanks and thanks everyone